The Web as a Broadcast/Publishing Hybrid

Matthew Gray

April 29, 1997

Introduction

Often, when the web is described is is contrasted to broadcast media and to publishing media. These comparisons to existing media forms typically are made to help people better understand potential applications and impact of the new media.

Broadcast Comparison

In many ways, the web has become used much like television or radio, as a way of delivering content to a large number of viewers. In some forms, such as RealAudio, StreamWorks, and other streaming forms, the comparison is inevitable, and fairly accurate.

For many other forms, such as news or entertainment web sites, the comparison is somewhat apt. The salient features of broadcast media tend to be a single high quality source disseminating information in one direction to a large number of viewers. Typically, there is advertising (as in the case of radio and network television), or a fee for receipt (as in the case of cable television). Additionally, the content is typically audiovisual. This clearly applies to many existing web sites and is a popular view.

Publishing Comparison

Another popular view of the web is to compare it to a large library of self published books, magazines, articles. The salient features of publishing tend to be similar to that of broadcasting, on close inspection. The information tends to be from a range or sources, and made available to a large number of readers. Typically the material is purchased, and sometimes contains advertising. Typically the content is largely textual, and somewhat graphical.

In terms of the web, this model fits even better from a content point of view. Despite many technologies enabling other forms of content, the vast majority of web content is still textual and grahical. The underlying economic structure, however is not terribly similar. Few for fee web sites exist outside of specific markets, and most free traditional publications are far more heavily advertising supported than the web currently is.

Differences in Broadcast and Publishing

In many ways it might appear that Broadcast and Publishing are the same, save for the form of their content and their economic support structure, but this is fairly clearly not the case.

Important differences lie in the media background and how they are perceived. Much of the difference is in the perceived accessibility. To the average individual, broadcast media are much more grandiose, and less accessible. Publishing on the other hand is seen as reachable. In this sense, an average person may consider (and execute) the publication of some content they generate, when they would never consider the same for a broadcast medium. This is substantially due to the perceived ease in creating the content.

Hybrid

In many ways, the web is a hybrid of these two media forms, though it does seem to be more of a publishing form. Some content, even simply text and images, could easily be considered broadcast, and yet the web is typically seen as rather approachable, as evidenced by the ubiquity of personal home pages.

This hybrid form has its own problems in terms of people identifying with one component and failing to realize that it is not excluslively a publishing or broadcast medium.

Conclusion

In conclusion, the web contains many of the facets of both a publishing medium and a broadcast medium. It does, however, not fit either of these models in a number of ways, partly by nature of its mixed forms (text, graphics, video, audio, application content). It is partly by the fact that it is less one directional than broadcast or publishing traditionally are. While a viewer or reader can write, call, fax or even email the producers of published or broadcast content, the response mechanism is not integrated into the viewing mechanism as it is in the case of the web.

Overall, the web is similar to broadcasting and publishing, but not sufficiently modeled by only these two comparisons.